1.the following data belong to co. for a period

material used $10000 (80%for DM)

labor used $50000 (70% for DL)

FOH cost (others) 5000

instruction:-compute

a. prime cost

b.conversion cost

c.production cost (mfg cost )

d.FOH cost

e.direct cost

f.indirect cost


1
Expert's answer
2021-08-17T10:23:00-0400

Solution:

a.). Prime cost = Direct Materials used + Direct Labor used

Direct Material used = "80\% \times 10,000 = 8,000"

Direct Labor used = "70\% \times 50,000 = 35,000"

Prime cost = "8,000 + 35,000 = 43,000"

 

b.). Conversion cost = Direct labor + Manufacturing Overheads

Conversion cost = 35,000 + 5,000 = 40,000

 

c.). Production cost (Mfg cost) = Direct Materials used + Direct Labor used + Manufacturing Overheads

Production cost (Mfg cost) = 8,000 + 35,000 + 5,000 = 48,000

 

d.). Factory Overhead Cost (FOH cost) = 5,000

 

e.). Direct costs = Direct material cost + Direct labor cost

Direct costs = 8,000 + 35,000 = 43,000

 

f.). Indirect costs = Manufacturing overheads

Indirect costs = 5,000